Your client is a 22-year old woman with a resting heart rate of 65. You perform a single-stage submaximal treadmill test at a Vo2 demand of 28 ml/kg/min, resulting in a heart rate of 132 bpm. What is her estimated Vo2max in ml/lk/min
Answer #1

Given

Age:22 years

Gender:Female

Maximum heart rate(MHR):132

Resting heart rate(RHR):65

formula:

Vo2 max=15.3*MHR/RHR

=15.3*132/65

=15.3*2

=31

Her estimated Vo2 will be 3ml/OK/min